OpenStudy (anonymous):

Solve the system by the substitution method. 2y-x=5 x^2 + y^2 - 25 = 0

OpenStudy (anonymous):

i know I have to do something with the first equation and then place it in the second but I am stuck there

OpenStudy (anonymous):

solve for x in the first equation... then use that x in the second equation.

OpenStudy (anonymous):

so that would make it x=-5+2y

OpenStudy (anonymous):

yes. now plug (-5+2y) in for x in the second equation.

OpenStudy (anonymous):

okay so now i have -5 + 2y + y^2 - 25 = 0

OpenStudy (anonymous):

no. (-5+2y)^2+y^2-25-0 you have to square the value of x

OpenStudy (anonymous):

okay now I am stuck

OpenStudy (anonymous):

(-5+2y)^2+y^2-25=0 expand the squared value

OpenStudy (anonymous):

the squares always get me

OpenStudy (anonymous):

okay I tried it and I got 5y^2 - 50

OpenStudy (anonymous):

\[(-5+2y)^2+y^2-25=0\] \[25-20y+4y^2+y^2-25=0\] \[5y^2-20y=0\] factor out the y \[y(5y-20)=0\] so y=0 or 5y-20=0 y=0 or y=4

OpenStudy (anonymous):

now use those two values of y and plug them back in where you had x=-5+2y you get x=-5 or x=(-5+8)=3

OpenStudy (anonymous):

but that don't match my choices I have (-5,0)

OpenStudy (anonymous):

yea your solutions are (-5,0) or (3,4)

OpenStudy (anonymous):

oh so x = -5 and 3 y= 0 and 4

OpenStudy (anonymous):

okay i see it now I forgot the plugging in of y to get x

OpenStudy (anonymous):

its all good. well as long either (-5,0) or (3,4) match your choices ur fine

OpenStudy (anonymous):

I was looking at what you did and I thought we found x first I was just confused

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome